what is the probability of rolling a 2,3, or 6 on a regular six sided die?
Answer #1

Since the number of choices or outcomes is six each “roll” of the die falls on either 1,2,3,4,5 or 6. That is, each roll has a probability of 1/6

We want to find, P( 2 or 3 or 6)

P(2 or 3 or 6)

= P(2) + P(3) + P(6)

= (1/6) + (1/6) + (1/6)

= 3/6

= 1/2

=> P(2 or 3 or 6) = 1/2 = 0.5

Probability = 1/2 = 0.5